OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 117r4 square inches. Each tile has an area of . The total number of tiles used can be represented by the expression below. Simplify the expression for the total number of tiles used. Show your work.

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

and each tile has an area of what?

OpenStudy (anonymous):

\[117r^{4} \over \sqrt{9}r ^{13} \]

OpenStudy (anonymous):

Total Area=Area of one tile * number of tiles

OpenStudy (anonymous):

Total Area=Area of one tile * number of tiles => 117r^4=sqrt(9r^13) *x => x=117r^4/sqrt(9r^13)=39/r^(5/2)